# Mixture-of-Experts with Gradient Conflict-Driven Subspace Topology Pruning for Emergent Modularity

### v1

@@ 4.3.3 Experiment III: The Necessity of Conflict (Pure Blind Baseline)
Figure 7 illustrates the routing distribution of the Pure Blind Baseline. Unlike CDSP, which achieves clear separation, this baseline exhibits severe Semantic Entanglement:
  • Resource Contention: Experts E0, E6, and E7 become "hotspot resources" shared indiscriminately by all tasks. The system gravitates towards a "Winner-Take-All" configuration dominated by these few experts.
@@
  • The "Pseudo-Balance" Trap: While the auxiliary loss prevents single-expert collapse, it forces an "egalitarian" distribution where distinct tasks are coerced into sharing experts to satisfy statistical uniformity.
Figure 7: Routing distribution of the Pure Blind Baseline at Epoch 9.
@@
Comparing Experiment I and III reveals the fundamental role of the conflict mechanism. Without the penalty for gradient interference, the model defaults to a "mixed sharing" strategy to minimize global loss, resulting in a chaotic overlap of symbolic and object features. The conflict game in CDSP is therefore identified as the essential force that breaks symmetry and drives the system toward modular disentanglement.
@@ 5 Conclusion
In this work, we challenged the prevailing design paradigm of Sparse Mixture-of-Experts, which relies on disjoint parameter storage and heuristic load-balancing constraints. We identified that such structural isolation, coupled with explicit task routing, leads to severe instruction overfitting and knowledge fragmentation. To address this, we introduced CDSP-MoE, a framework that grounds expert specialization in the physical interaction of gradients within a shared subspace.
Our contributions are threefold. First, by replacing static routing tables with a dynamic, conflict-driven pruning mechanism, we demonstrated that modularity can emerge purely from the minimization of physical interference. Second, our experiments revealed a critical trade-off: CDSP-MoE initially lags behind the baseline in classification accuracy during early training phases. We interpret this as an evolutionary tax—while the baseline rapidly minimizes loss by memorizing simple ID-to-expert mappings ("shortcut learning"), CDSP-MoE must invest compute cycles to resolve gradient conflicts and restructure its topology. However, this initial cost yields a significant structural dividend: unlike the baseline which collapses under blind inference, CDSP-MoE evolves robust, content-aware routing pathways that disentangle symbolic concepts from object features without human supervision. Third, the comparison with the Pure Blind Baseline highlighted that auxiliary load-balancing losses alone are insufficient for semantic decoupling; the adversarial signal provided by the gradient conflict game is essential for breaking symmetry.
@@
While CDSP-MoE offers a principled path toward autonomous modularity, it incurs a computational cost. Calculating the gradient conflict matrix requires storing lagged gradients, which increases memory overhead compared to standard forward-only gating. Additionally, our current validation is limited to vision classification tasks. Future work will focus on scaling this architecture to Large Language Models, investigating whether conflict-driven pruning can spontaneously separate distinct reasoning capabilities (e.g., coding vs. creative writing) and mitigate catastrophic forgetting in continuous learning scenarios. We believe that grounding neural architecture search in physical gradient dynamics represents a promising step toward interpretable and self-organizing artificial intelligence.
Appendix A Appendix
A.1 Detailed Hyperparameters
To ensure reproducibility, we list the detailed hyperparameters used for the CDSP-MoE experiments in Table A1.
Table A1: Hyperparameter settings for the Heterogeneous Multi-Task Experiment.
@@ A.1 Detailed Hyperparameters
Topology Learning Rate ($\eta_{topo}$) $5\times 10^{-2}$ ($10\times$)

### v3

@@ 4.3.3 Experiment III: The Necessity of Conflict (Pure Blind Baseline)
Figure 3(c) illustrates the routing distribution of the Pure Blind Baseline. Unlike CDSP, which achieves clear separation, this baseline exhibits severe Semantic Entanglement:
  • Resource Contention: Experts E0, E6, and E7 become ”hotspot resources” shared indiscriminately by all tasks. The system gravitates towards a ”Winner-Take-All” configuration dominated by these few experts.
@@
  • The ”Pseudo-Balance” Trap: While the auxiliary loss prevents single-expert collapse, it forces an ”egalitarian” distribution where distinct tasks are coerced into sharing experts to satisfy statistical uniformity.
@@
Comparing Experiment I and III reveals the fundamental role of the conflict mechanism. Without the penalty for gradient interference, the model defaults to a ”mixed sharing” strategy to minimize global loss, resulting in a chaotic overlap of symbolic and object features. The conflict game in CDSP is therefore identified as the essential force that breaks symmetry and drives the system toward modular disentanglement.
@@ 5 Conclusion
In this work, we challenged the prevailing design paradigm of Sparse Mixture-of-Experts, which relies on disjoint parameter storage and heuristic load-balancing constraints. We identified that such structural isolation, coupled with explicit task routing, leads to severe instruction overfitting and knowledge fragmentation. To counter this, we introduced CDSP-MoE, a framework that grounds expert specialization in the physical interaction of gradients within a shared subspace.
Our contributions are threefold. First, by replacing static routing tables with a dynamic, conflict-driven pruning mechanism, we demonstrated that modularity can emerge purely from the minimization of physical interference. Second, our experiments revealed a critical trade-off: CDSP-MoE initially lags behind the baseline in classification accuracy during early training phases. We interpret this as an evolutionary tax—while the baseline rapidly minimizes loss by memorizing simple ID-to-expert mappings (”shortcut learning”), CDSP-MoE must invest compute cycles to resolve gradient conflicts and restructure its topology. However, this initial cost yields a significant structural dividend: unlike the baseline which collapses under blind inference, CDSP-MoE evolves robust, content-aware routing pathways that disentangle symbolic concepts from object features without human supervision. Third, the comparison with the Pure Blind Baseline highlighted that auxiliary load-balancing losses alone are insufficient for semantic decoupling; the adversarial signal provided by the gradient conflict game is essential for breaking symmetry.
@@
While CDSP-MoE offers a principled path toward autonomous modularity, it incurs a computational cost. Calculating the gradient conflict matrix requires storing lagged gradients, which increases memory overhead compared to standard forward-only gating. However, this is an engineering rather than theoretical bottleneck; in large-scale models, this overhead can be mitigated by low-rank gradient projection or sparse conflict sampling (computing conflicts only for top-$k$ active experts), making the approach scalable. Additionally, our current validation is limited to vision classification tasks. Future work will focus on scaling this architecture to Large Language Models, investigating whether conflict-driven pruning can spontaneously separate distinct reasoning capabilities (e.g., coding vs. creative writing) and resist catastrophic forgetting in continuous learning scenarios. We believe that grounding neural architecture search in physical gradient dynamics represents a promising step toward interpretable and self-organizing artificial intelligence.
Appendix
This appendix provides supplementary materials supporting the main paper’s findings. The content is organized as follows:
  • Appendix A details the exact hyperparameters and architectural configurations used in the experiments.
  • Appendix B presents extended visualizations of the topology matrix evolution and routing distributions.
  • Appendix C and D verify the training stability and analyze the structural convergence dynamics.
  • Appendix E provides the rigorous theoretical proofs deriving the inevitability of modular emergence from gradient conflict dynamics.
Appendix A Detailed Hyperparameters
Table A1: Hyperparameter settings.
@@ Appendix A Detailed Hyperparameters
Topology Learning Rate ($\eta_{topo}$) $5\times 10^{-2}$